# LightTact: A Visual–Tactile Fingertip Sensor for Deformation-Independent Contact Sensing

### v1

@@ III-C Sensor Fabrication
Fabricating the two gel layers of LightTact requires several auxiliary molds and temporary inserts to form accurate internal boundaries, as illustrated in Fig. 4.
Transparent gel. The transparent gel is cast from Silicones Inc® XP-565 silicone (mixing ratio 10:1, shore A27), chosen for its high clarity and stable mechanical properties. The acrylic window, laser-cut from $1.5\text{mm}$ acrylic sheet, is installed inside the 3D-printed black PLA shell. To create a sealed cavity for casting the transparent silicone, the LED mount geometry is temporarily modified by filling its original LED recess with a solid insert (labeled 0 in Fig. 4(a)) so that no silicone leaks into the LED compartment. Two additional 3D-printed molds (labeled 1 and 2) are stacked around the shell to reserve the perimeter region occupied by the future black gel.
Before pouring the transparent silicone, a high-absorption black pigment (Black 2.0) is brushed onto the relevant inner surfaces of the shell (labeled A in Fig. 4(a)) to more closely approximate the idealized absorbing boundaries assumed in Section III-A. To ensure strong bonding between the acrylic window and the cured silicone, a thin layer of silicone adhesive (Smooth-On® Sil-Poxy) is brushed onto the acrylic face that will contact the gel (labeled B in Fig. 4(a)). After preparing these surfaces, degassed XP-565 silicone is poured into the cavity until it completely fills the interior of the shell. It takes about 4 hours for the silicone to fully cure on a heated 3D-printer bed maintained at $50\,^{\circ}\mathrm{C}$.
Black gel. The black gel is made from the same XP-565 silicone mixed with black pigment. After the transparent silicone cures, the upper mold (2) is removed and replaced by a second top mold (labeled 3 in Fig. 4(b)), which defines the perimeter volume allocated the black gel. Because this cavity is narrow and difficult to fill uniformly from above, we place an additional acrylic mold (labeled 4) on top of mold 3. The mixed black silicone is poured through the opening in mold 4 and flows into the perimeter cavity underneath it. This approach keeps the acrylic mold firmly pressed against the touching surface and ensures that the black gel cures to a uniform height around the transparent gel. After curing under the same heated-bed conditions, all molds are removed and excess gel around the fill opening is trimmed.
As indicated by label C in Fig. 4(b), the shell includes a small raised step along its perimeter. This feature creates the final three-layer edge profile (“shell–black gel–transparent gel”). The reason for this geometry is that mold 4 applies slight pressure on the transparent gel during the black-gel casting step. Once the molds are removed, the transparent gel relaxes back to its undeformed shape, which lifts the cured black gel upward by a small amount and leaves a thin air gap between the underside of the black gel and the top surface of the shell. Without the raised step on the shell perimeter, this air gap would open a direct lateral path for external light to enter the sensor and reach the camera. The raised perimeter therefore acts as a light-blocking barrier, preserving the “black surface” assumption required by our optical design.
@@ III-D Contact Segmentation
Camera operation and sensor calibration. The camera is set with auto-exposure disabled and uses a fixed exposure time of $20\penalty 10000\ \text{ms}$ during operation. This is crucial because auto-exposure would brighten the predominantly dark background of LightTact, amplifying sensor noise and potentially saturating bright contact regions. For robotic use, we additionally calibrate image pixels to their spatial positions on the touching surface. Following 9DTact [lin20239dtact], we design a calibration tool with a $5\times 5$ array of cylindrical bumps at a spacing of $3\penalty 10000\ \text{mm}$. As shown in Fig. 5(a), LightTact captures the imprint pattern when the tool is pressed onto the touching surface. We segment the imprints by simple thresholding and extract their center pixels (Fig. 5(b)). We then select the four outermost corner imprints in green to estimate the average pixel spacing, and compute the target rectified locations of all centers on an evenly spaced grid. Finally, using the correspondences between the detected centers and their rectified grid locations, we compute a warping map to rectify and crop each raw image to a top-down view (Fig. 5(c)).
Segmentation algorithm. The optical design of LightTact naturally produces a high-contrast representation of contact: non-contact regions remain near-black, while true contact results in localized brightness increases. Building on this property, our segmentation algorithm adopts a simple yet highly robust frame-differencing strategy. Before operation, LightTact records $N$ no-contact frames and computes their pixel-wise average to form a stable reference image $I_{\mathrm{ref}}$. For each incoming frame $I$, we compute a difference image $I_{\mathrm{diff}}=I_{\mathrm{raw}}-I_{\mathrm{ref}}$, in which contact pixels appear as positive brightness changes. To reliably distinguish physical contact from sensor noise, we introduce a multi-condition brightness consistency check. A pixel at location $(x,y)$ is classified as contact if any of the following four conditions is satisfied: (1) the mean RGB increase exceeds a threshold $t_{0}$; (2) at least one color channel exceeds threshold $t_{1}$; (3) at least two color channels exceed threshold $t_{2}$; or (4) all three color channels exceed threshold $t_{3}$. These complementary conditions enable the algorithm to detect both prominent brightness changes and more subtle, consistent increases produced by contact. In our implementation, we use $N=10$ and thresholds $\{t_{0},t_{1},t_{2},t_{3}\}=\{25,20,30,40\}$.
@@ IV-A Light Suppression at Non-Contact Regions
We begin by evaluating whether LightTact achieves the intended light-suppression behavior under varying levels of internal illumination and external ambient brightness. We measure the brightness at the touching surface using a digital light meter. For each condition, we record (1) raw images without contact (non-contact images) and (2) raw images with a cylindrical object touching the sensing surface (contact images). We report the mean and standard deviation of the non-contact images in Table II, along with representative raw and segmented results.
@@ IV-B Deformation-Independent Contact Segmentation
Extremely light contact sensing. Figure 6 evaluates LightTact’s ability to detect contact from liquids (green juice, milk), semi-liquids (toothpaste), and ultra-soft materials (cotton, sponge, tofu, noodle, beef, fingertip, palm) under bright indoor lighting. All objects are placed or lightly touched onto the sensor without pressing, producing minimal or no macroscopic surface deformation. Despite this, LightTact reliably “sees” the contact region directly, and its optical light-suppression mechanism yields clean, pixel-level contact segmentation even for objects with complex appearance. In contrast, 9DTact [lin20239dtact] and GelSight [yuan2017gelsight] fail to provide even binary contact detection under such light interactions, as their sensing fundamentally depends on measurable deformation. Moreover, LightTact can sense contact even under effectively zero applied force as demonstrated by the thin film hanging upside down on the downward-facing LightTact sensor in Fig. 1, indicating that LightTact does not require a minimum force threshold for detection.
@@
Figure 7: LightTact reliably senses both light and firm contact from rigid objects while preserving their visual appearance.
Beyond liquids and soft materials, LightTact also provides robust visual–tactile sensing for rigid objects, as shown in Fig. 7 (textured joystick surface, complex-pattern cylinder, beads, cube, AirPods). Deformation-based VBTSs sense only indentation and become ineffective under light contact, as illustrated by the bead example. In contrast, LightTact consistently produces clean segmentation and natural visual appearance.
Together, these results demonstrate that LightTact reliably detects contact and performs accurate pixel-level segmentation across liquids, semi-liquids, ultra-soft materials, and rigid objects. This level of robustness to material properties, surface appearance, and environmental lighting is not achievable with TIR-based VBTSs, and the multimodal richness is far beyond what deformation-based tactile sensors can provide.
@@ IV-C1 Light contact sensing of liquids, semi-liquids, and soft materials
When a robot approaches an object, the wrist camera view is often occluded, making tactile sensing essential for safe interaction. Materials without rigid boundaries (e.g., water and creams) and extremely soft or thin materials (e.g., food film) are particularly challenging for force-based and deformation-based tactile sensors because their contacts generate unmeasurable forces or microscopic deformation. In contrast, LightTact can directly detect such light contacts through its deformation-independent optical principle, enabling to unlock such light-contact interactions for robots.
@@ IV-C2 Multimodal sensing and VLM-guided fine-grained manipulation
In contrast, the baseline sensors do not provide interpretable appearance cues and therefore cannot support this VLM-guided sorting pipeline. We also compare against wrist-camera images: the VLM fails consistently because the resistor occupies a small fraction of the image and the background is cluttered. Overall, these results demonstrate that LightTact can provide VLM-readable inputs without fine-tuning, enabling robots to obtain high-level task decisions directly through prompting.
V Conclusion
We presented LightTact, a compact visual-tactile fingertip sensor that enables deformation-independent, optics-based contact sensing. The core idea is an ambient-blocking optical configuration that suppresses both external light and internal illumination at non-contact regions, while transmitting only the diffuse light generated at true contacts. This yields high-contrast raw images with near-black non-contact pixels and appearance-preserving contact pixels, making pixel-level contact segmentation simple and robust. Across systematic evaluations, LightTact maintains reliable contact segmentation over diverse materials (liquids, semi-liquids, ultra-soft materials, and rigid objects) and under strong ambient lighting. When integrated on a robotic arm, LightTact enables manipulation behaviors driven by extremely light contact, including water spreading, facial-cream dipping, and thin-film interaction, where force-based sensing and deformation-based VBTSs fail. Moreover, LightTact provides spatially aligned visual-tactile images that can be directly prompted to vision-language models for fine-grained reasoning, demonstrated by resistor value inference for robotic sorting.
@@ Acknowledgments
We would like to thank Yuxiang Yang for his insightful discussions on sensor design and experiments throughout the project.
@@
V-A Camera Placement under $\theta_{tv}<2\theta_{c}$
@@
V-B Design of Sensor Components
@@
V-C Optical Analysis with Wedge Shape ($\theta_{s}<\frac{\pi}{2}$)
@@
As $\theta_{s}$ decreases, the slanted internal shell surface becomes more likely to reflect ambient light (both external and internal) into directions that can enter the camera through non-contact regions, leading to a gradual increase in non-contact intensity. Nevertheless, even at $\theta_{s}=30^{\circ}$, the non-contact images remain close to black (mean gray value <4.0<4.0), indicating that the light-suppression property is largely preserved under moderate wedge shaping. If even darker non-contact images are desired, the slanted surface can be further improved by enhancing its absorptivity (e.g., applying higher-absorption black coating) or by adding micro-textures/light-trap patterns to reduce specular reflections.
V-D Large-Area Contact Sensing
@@
V-E Details for Resistor Sorting

### v2

@@ III-C Sensor Fabrication
Most components of LightTact are straightforward to fabricate: the shell and mounts are 3D-printed with black PLA, the acrylic window is laser-cut from $1.5\penalty 10000\ \text{mm}$ acrylic sheet, the LED board is made by FPCB services, and the camera is off-the-shelf. The only nontrivial steps are casting the two gel layers, which require auxiliary molds and temporary inserts to define accurate internal boundaries (Fig. 4).
Transparent gel. We choose Silicones Inc® XP-565 silicone (cured hardness Shore A27) to make the transparent gel. The acrylic window is installed in the shell to form a sealed container for casting. To prevent silicone leakage into the LED compartment, we temporarily fill the LED recess with a solid insert (mold 0 in Fig. 4(a)). Two additional 3D-printed molds (labeled 1 and 2) are stacked around the shell to reserve the perimeter volume occupied by the future black gel.
Before pouring, we brush a high-absorption black pigment onto the relevant inner shell surfaces (region A in Fig. 4(a)) to better approximate the light-absorbing boundary condition. To ensure strong bonding at the gel-acrylic interface, we apply a thin layer of silicone adhesive (Smooth-On® Sil-Poxy) to the acrylic face labeled B. We then degas and pour the XP-565 mixture until the cavity is fully filled, and cure it for $\sim$4 hours on a heated 3D-printer bed at $50^{\circ}\mathrm{C}$.
Black gel. The black gel is made from the same XP-565 silicone mixed with black pigment. After the transparent gel cures, we remove mold 2 and replace it with a second top mold (mold 3 in Fig. 4(b)) that defines the perimeter cavity for the black gel. Because this cavity is narrow and difficult to fill uniformly from above, we place an additional acrylic mold (labeled 4) on top of mold 3 and pour the mixture through the opening in mold 4, allowing it to flow uniformly around the perimeter. This “cap” mold also keeps firmly pressed against the touching surface, ensuring a consistent cured height. After curing under the same heated-bed conditions, we remove the molds and trim any excess gel around the fill opening.
As indicated by label C in Fig. 4(b), the shell includes a small raised perimeter step that creates the final three-layer edge profile (“shell – black gel – transparent gel”). This step is necessary because mold 4 slightly compresses the transparent gel during black-gel casting; after demolding, the transparent gel elastically recovers and can leave a thin lateral air gap beneath the black gel. Without the raised step, this gap would form a direct light path into the sensor. The raised perimeter therefore acts as a light-blocking barrier, preserving the dark-boundary assumption required by our optical design.
@@ III-D Contact Segmentation
Camera operation and sensor calibration. We disable auto-exposure and operate the camera at a fixed exposure time of $20\penalty 10000\ \text{ms}$. This is crucial because auto-exposure would brighten LightTact’s predominantly dark background, amplifying sensor noise and potentially saturating bright contact regions. For robotic use, we further calibrate a mapping from image pixels to physical locations on the touching surface. Following 9DTact [lin20239dtact], we use a calibration tool with a $5\times 5$ array of cylindrical bumps at $3\penalty 10000\ \text{mm}$ spacing. We press the tool onto the surface and capture the resulting imprint pattern (Fig. 5(a)). We segment the imprints via thresholding and extract their center pixels (Fig. 5(b)). We then use the four outermost corner imprints to estimate the average pixel spacing and define the target rectified grid. Finally, we compute a warping map from the detected centers to the rectified grid, and apply it to rectify and crop each raw image into a top-down view (Fig. 5(c)).
Segmentation algorithm. The optical design of LightTact naturally produces a high-contrast representation of contact: non-contact regions remain near-black, while true contact results in localized brightness increases. Building on this property, our segmentation algorithm adopts a simple yet highly robust frame-differencing strategy. Before operation, LightTact records $N$ no-contact frames and computes their pixel-wise average as a reference image $I_{\mathrm{ref}}$. Given an incoming frame $I$, we compute a difference image $I_{\mathrm{diff}}=I_{\mathrm{raw}}-I_{\mathrm{ref}}$, in which contact pixels appear as positive brightness changes.
To separate true contact from sensor noise, we apply a multi-condition brightness consistency test. A pixel at location $(x,y)$ is classified as contact if any of the following holds: (1) the mean RGB increase exceeds $t_{0}$; (2) at least one channel exceeds $t_{1}$; (3) at least two channels exceed $t_{2}$; or (4) all three channels exceed $t_{3}$. These complementary criteria capture both strong changes and weaker but consistent increases produced by contact. In our implementation, we use $N=10$ and ${t_{0},t_{1},t_{2},t_{3}}={25,20,30,40}$.
@@ IV-A Light Suppression at Non-Contact Regions
We first evaluate whether LightTact achieves the intended light-suppression behavior under varying levels of internal illumination and external ambient brightness. We measure the brightness at the touching surface using a digital light meter. For each condition, we record (1) raw images without contact (non-contact images) and (2) raw images with a cylindrical object touching the sensing surface (contact images). We report the mean and standard deviation of the non-contact images in Table II, along with representative raw and segmented results.
@@ IV-B Deformation-Independent Contact Segmentation
Extremely delicate contact sensing. Figure 6 evaluates LightTact’s ability to segment contact for liquids (green juice, milk), semi-liquids (toothpaste), and ultra-soft materials (cotton, sponge, tofu, noodle, beef, fingertip, palm) under bright indoor illumination. Across all trials, objects are placed or gently touched onto the sensor without intentional pressing, resulting in minimal or no macroscopic surface deformation. Despite the absence of indentation cues, LightTact directly identifies the contact region from light intensity changes in the captured image, as its optical light-suppression mechanism produces clean, pixel-level segmentation even for visually complex materials. By contrast, deformation-based tactile sensors, including 9DTact [lin20239dtact], GelSight [yuan2017gelsight], and DelTact [zhang2022deltact], do not yield reliable contact detection in this near-zero-force regime, since their measurements fundamentally rely on measurable deformation. Moreover, LightTact can detect contact under effectively zero applied force, as shown by the thin film hanging upside down on the downward-facing LightTact in Fig. 1; this further indicates that LightTact does not require a minimum force threshold to trigger contact perception.
@@
Figure 7: LightTact reliably senses both light and firm contact from rigid objects while preserving their visual appearance. In contrast, deformation-based tactile sensors fail under light contact, as the bead example illustrated.
Beyond liquids and soft materials, LightTact also provides robust visual–tactile sensing for rigid objects, as shown in Fig. 7 (textured joystick surface, complex-pattern cylinder, beads, cube, AirPods). Across both gentle placement and firmer pressing, LightTact consistently produces accurate contact segmentation while preserving natural visual appearance in the raw image. In contrast, deformation-based VBTSs become unreliable with rigid bodies when contact does not induce sufficient deformation (e.g., the bead case). Overall, these results show that LightTact performs accurate, pixel-level contact segmentation across liquids, semi-liquids, ultra-soft materials, and rigid objects. Such robust coverage across materials, appearances, and lighting is difficult to achieve with both TIR-based and deformation-based VBTSs.
@@ IV-C1 Light contact sensing of liquids, semi-liquids, and soft materials
When a robot approaches an object for grasping or manipulation, the wrist-camera view is often occluded by the end-effector and the object, making tactile sensing essential for safe, closed-loop interaction. However, most VBTSs rely on contact-induced deformation and thus provide weak or unreliable signals when contact produces little-to-no indentation. In these regimes, LightTact remains effective: it detects contact at the gripper via a deformation-independent optical principle.
@@ IV-C2 Multimodal sensing and VLM-guided fine-grained manipulation
In contrast, the baseline sensors do not provide interpretable appearance cues; therefore cannot support this VLM-guided sorting pipeline. We also compare against wrist-camera images: the VLM fails consistently because the resistor occupies a small fraction of the image and the background is cluttered. Overall, these results demonstrate that LightTact can provide VLM-readable inputs without fine-tuning, enabling robots to obtain high-level task decisions directly through prompting.
V Ablation Study for Sensor Design
We evaluate ablation studies for several key design and fabrication choices introduced in Section III, with full results reported in Appendix VI-D.
  • Black paint on the shell inner walls
  • Acrylic mold for casting the black gel
  • Edge indentation on the shell outer walls
VI Conclusion
We presented LightTact, a compact visual-tactile fingertip sensor that enables deformation-independent, optics-based contact sensing. The core idea is an ambient-blocking optical configuration that suppresses both external light and internal illumination at non-contact regions, while transmitting only the scattered light generated at true contacts. This yields high-contrast raw images with near-black non-contact pixels and appearance-preserving contact pixels, making pixel-level contact segmentation simple and robust. Across systematic evaluations, LightTact maintains reliable contact segmentation over diverse materials (liquids, semi-liquids, ultra-soft materials, and rigid objects) and under strong ambient lighting. When integrated on a robotic arm, LightTact enables manipulation behaviors driven by extremely light contact, including water spreading, facial-cream dipping, and thin-film interaction, where force-based sensing and deformation-based VBTSs fail. This is promising for applications such as delicate handling of cosmetics, food, and biomedical samples, operating on wet surfaces, and enabling safe human-robot interaction. Moreover, LightTact provides spatially aligned visual-tactile images that can be directly prompted to vision-language models for fine-grained reasoning, demonstrated by resistor value inference for robotic sorting.
@@ Acknowledgments
We would like to thank Yuxiang Yang for his insightful discussions throughout the project.
@@
VI-A Camera Placement under $\theta_{tv}<2\theta_{c}$
@@
VI-B Design of Sensor Components
@@
VI-C Optical Analysis with Wedge Shape ($\theta_{s}<\frac{\pi}{2}$)
@@
As $\theta_{s}$ decreases, the slanted internal shell surface becomes more likely to reflect ambient light (both external and internal) into directions that can enter the camera through non-contact regions, leading to a gradual increase in non-contact intensity. Nevertheless, even at $\theta_{s}=30^{\circ}$, the non-contact images remain close to black (mean gray value <4.0<4.0), indicating that the light-suppression property is largely preserved under moderate wedge shaping. If even darker non-contact images are desired, the slanted surface can be further improved by applying higher-absorption black coating or by adding micro-textures/light-trap patterns to reduce specular reflections.
VI-D Ablation Study For Sensor Design
Acrylic mold for casting the black gel
Without the acrylic mold, the narrow perimeter cavity makes it difficult to control the volume of black silicone, often resulting in either overflow or underfill. Overflow contaminates the sensing surface and permanently blocks contact visibility in the covered regions, while underfill exposes side surfaces and opens a direct leakage path for ambient light to reach the camera (Fig. 14 (a)).
Figure 14: Underfill and the missing perimeter step of the black gel both allow external light to leak into the sensor.
Edge indentation on the shell outer walls
As shown in Fig. 14 (b), removing the perimeter step (via the edge indentation design) prevents the black gel from forming a light-blocking lip at the shell boundary. This leaves a thin gap between the shell and the gel, which again creates a leakage path for ambient light into the sensor and raises the non-contact background.
Black paint on the shell inner walls
As shown in Fig. 15, brushing high-absorption black paint on the inner walls significantly darkens the internal surfaces. In practice, the unpainted shell reflects more stray light toward the camera as ambient brightness increases, leading to a brighter non-contact background and reduced contrast.
Figure 15: Applying high-absorption black paint substantially darkens the shell inner walls and reduces stray reflections.
VI-E Large-Area Contact Sensing
@@
VI-F Baseline Vision-Based Tactile Sensors
Our baseline VBTSs include GelSight, DelTact and 9DTact, as shown in Fig. 17.
Figure 17: Three baseline sensors and LightTact.
VI-G Details for Resistor Sorting